OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope (3, –10); m = –0.8. Just need a second opinion

OpenStudy (anonymous):

y-y1=m(x+x1) y --10= -0.83(x + 3) y + 10= -0.83(x +3)